# Raw Data Matters: Enhancing Prompt Tuning by Internal Augmentation on Vision-Language Models

### v1

@@ Introduction
Bridging images and text, Vision-Language Models (VLMs, represented by CLIP (Radford et al. 2021)) pre-trained on hyper-scale image-text pairs have demonstrated impressive cross-modal alignment and fusion capabilities (Li et al. 2023; Alayrac et al. 2022). To efficiently adapt pretrained VLMs for classification-related downstream tasks, CLIP-based prompt tuning is proposed as a parameter-efficient fine-tuning (PEFT) approach (Xing et al. 2024). Freezing all parameters in CLIP encoders, prompt tuning introduces a set of lightweight learnable prompt vectors that replace hard prompts in raw inputs (Zhou et al. 2022b, a). As self-adaptive queries, prompt vectors guide the pretrained CLIP toward the target distribution through efficient fine-tuning.
@@
Given the rareness of annotated image-text pairs in target downstream tasks, extant prompt-tuning strategies (Khattak et al. 2023a; Yao, Zhang, and Xu 2023; Khattak et al. 2023b) are generally carried out under few-shot settings. Therefore, intuitively, introducing more relevant data to the tuning process can enhance recognition performance on target categories (base classes). Empirical evidence from numerous studies (Zhou et al. 2022b; Khattak et al. 2023a, b) supports this conclusion. As illustrated in Fig. 1(a), with the increase of provided image-text pairs (shots), consistent improvement on fine-tuning performance can be observed on multiple backbone models. This indicates that the prompt tuning process is highly sensitive to the scale of data.
@@
Figure 1: (a) Impact of the amount of data (shots) per class on performance in prompt tuning models (Khattak et al. 2023a; Yao, Zhang, and Xu 2023; Khattak et al. 2023b); (b) Visualization of bad cases in existing self-supervised augmentation approaches. In comparison, AugPT obtains quality improvement as in (c) by filtering noisy samples, surpassing the current state-of-the-art PromptKD (Li et al. 2024a) as backbone model in (d) base-to-new tasks in both full and few-shot scenarios over 11 datasets.
@@
Motivated by this characteristic, numerous approaches explore optimizing prompt tuning at the data level. Current studies primarily focus on introducing external knowledge resources to boost base-class performance and generalization on out-of-distribution categories (new classes), containing pre-built hard prompt templates (Yao, Zhang, and Xu 2023; Yu et al. 2023) or structured knowledge bases (Kan et al. 2023; Zhang et al. 2024b), and knowledge generated by Large Language Models (LLMs) (Wang et al. 2024; Tian et al. 2024). Despite the effectiveness, these strategies require additional data collection and organization costs. In addition, reliance on task-specific external knowledge also limits the generality and transferability for fine-tuning the models. Latest research (Li et al. 2024a; Wu et al. 2024; Li et al. 2025a) further proposes methods for data amplification utilizing internal knowledge of the model. Typically guided by knowledge distillation, these approaches first fine-tune a prompt tuning backbone under few-shot settings (Khattak et al. 2023b) as teacher model, then guide it to execute knowledge completion for available unlabeled images for further fine-tuning a student model. Although distillation-based method achieve state-of-the-art (SOTA) in both base-class accuracy and new-class generalization, there are still notable limitations: (1) Performance gains still depend on the expansion of unlabeled image library, which also requires extra data collection costs; (2) Extracted image features are only applied for aligning teacher and student, signifying that raw images are not adequately exploited.
@@
It is worth noting that traditional self-supervised augmentation strategies (Cubuk et al. 2020; DeVries and Taylor 2017; Mao et al. 2022) face a key limitation: they lack an error correction mechanism, which may introduce noisy samples that discard key visual features, as illustrated in Fig. 1(b). Although this limitation can be addressed with additional supervision (Sohn et al. 2020), to maintain parameter efficiency of prompt tuning, we consider strategies that do not require external supervision. As an effective solution, AugPT introduces a novel Consensus Test filtering gate, leveraging a previously overlooked property of pre-tuned prompt tuning models: logit-level consistency reflects semantic similarity across images (details are in Method section). Concretely, we reuse the frozen teacher model (Li et al. 2024a) to perform online inference over both raw and augmented images, and adopt Top-1 frequency consistency as consensus to build the view filtering strategy, ensuring that core semantic features of raw image are preserved in augmented views (Fig. 1(c)). Without the need of external modules, this design injects a self-corrective mechanism into augmentation, thereby enhancing cross-modal alignment.
Experiments across 11 datasets reveal that AugPT exceeds PromptKD (Li et al. 2024a) (current SOTA) as backbone model on base-class performance, new-class generalization, and cross-dataset transfer, without using additional knowledge or external data. Our main contributions are:
@@
Prompt Tuning in VLMs. Although CLIP (Radford et al. 2021)-based VLMs has remarkable performance on cross-modal tasks, the Transformer-based encoders (Dosovitskiy et al. 2020) with large parameters make it challenging on full fine-tuning for adapting downstream tasks. Therefore, prompt tuning is proposed as a parameter-efficient strategy (Xing et al. 2024; Han et al. 2024). It does not rely on hard prompts (e.g., “A photo of a [CLASS]”), but introduces a set of learnable vectors as substitutes. During fine-tuning, all parameters of encoders are frozen, and the prompt vectors used as queries are continuously optimized to fit the distribution of target tasks. Existing research explores various formats of prompts, containing text re-encoding (Zhou et al. 2022b; Zhu et al. 2023; Tian et al. 2024), concatenation with image features (Jia et al. 2022), and joint visual-text encoding (Khattak et al. 2023a, b; Li et al. 2025b; Guo and Gu 2025). Since CLIP-based prompt tuning is typically conducted under few-shot scenarios, applying data amplification strategies like AugPT to enhance performance is intuitive, which is supported by the results in Fig. 1(a).
[…]
Data-Driven Optimization for Prompt Tuning. Given the limitation of image-text pairs for fine-tuning in CLIP-based prompt tuning, a variety of studies discover data-level optimization strategies. Mainstream approaches consider introducing external knowledge to expand data or constrain the optimizing flow, including elaborately designed hard prompt templates as supervision (Yao, Zhang, and Xu 2023; Khattak et al. 2023b), or structured databases for appending additional knowledge (Kan et al. 2023; Zhang et al. 2024b). More recently, LLMs are employed as external sources to generate descriptions (Pratt et al. 2023; Wang et al. 2024; Khattak et al. 2025), auxiliary attributes (Tian et al. 2024; Li et al. 2024b; Zhu et al. 2024), or guide fine-tuning (Zheng et al. 2024). Since the above methods demand extra constructing cost for external knowledge bases, the latest works (Mistretta et al. 2024; Li et al. 2024a) propose distillation-based paradigms using internal knowledge + external data. Specifically, a teacher model is first fine-tuned on few-shot image-text pairs and then guides a student model by knowledge completion to external unlabeled images. However, further optimizing such models still requires scaling up external data pool, signifying additional collection overhead. In contrast, AugPT proposes an optimized distillation-based strategy that relies only on internal knowledge + internal augmentation based on existing raw data.
@@
The framework of AugPT is illustrated in Fig. 2. AugPT adopts the infrastructure of an up-to-date distillation-based prompt tuning model (taking SOTA PromptKD (Li et al. 2024a) as a paradigm) as its backbone. Existing unlabeled images from the backbone are sequentially processed through an Adaptive Self-supervised Augmentation module and a Consensus-based Filtering Gate. The pipeline is then optimized for further exploiting latent features within the raw data during fine-tuning.
@@
[…]
@@
To avoid introducing additional learnable layers or external knowledge, thereby minimizing the cost of data collection and computational overhead, AugPT adopts parameter-free image augmentation based on self-supervised strategy. Replacing the naive Random Resized Crop (Mao et al. 2022) applied in standard prompt tuning, based on the RandAugment (Cubuk et al. 2020) methodology, AugPT performs $N$ augmentations on each raw unlabeled image $I^{\prime}$. For each augmentation, the model samples an operation from a pre-designed set with $Q$ candidates of mainstream transformation policies $\mathcal{Y}=\{\gamma_{1},\dots,\gamma_{Q}\}$ (details are in Supplementary Material A.3). Guided by a searching space defined by step size $S$ and amplitude $A$, augmentation operations $\gamma^{(s)}\sim\text{Uniform}(\mathcal{Y})$ are independently and identically sampled at each step $s\in[S]$ to form a policy group $\mathcal{G}$:
@@
Based on the filtered image set $\mathcal{I}_{\text{acc}}$ and the consensus category $\theta^{\star}$, AugPT executes further optimization following the fine-tuning workflow of distillation-based prompt tuning backbone models (Li et al. 2024a). Specifically, under the constraint of $\theta^{\star}$, AugPT samples the teacher logits ^$\hat{\mathcal{\ell}}^{\text{tea}}_{p}$ corresponding to each accepted image $\hat{I}_{p}\in\mathcal{I}_{\text{acc}}$ to organize a set $\mathcal{T}^{\text{tea}}_{\text{acc}}$ in a one-to-one correspondence manner:
@@
Next, AugPT constructs a student model identical to the backbone (which a frozen CLIP-ViT-B/16 image encoder $f^{\text{stu}}(\cdot)$ and learnable student image prompts $\boldsymbol{P}_{v}^{\text{stu}}$). Similarly, KL divergence is utilized to align the distributions between teacher logits ^$\hat{\mathcal{\ell}}^{\text{tea}}_{p}$ from the filtered set and student logits ^$\hat{\mathcal{\ell}}^{\text{stu}}_{p}=sim[g^{\text{tea}}({\boldsymbol{P}_{t}^{\text{tea}}}),\text{Proj}(f^{\text{stu}}({\boldsymbol{P}_{v}^{\text{stu}}},\hat{I_{p}}))]$, which are obtained by interacting the learnable student visual inputs $(\boldsymbol{P}_{v}^{\text{stu}},\hat{I}_{p})$ with the frozen teacher text prompts $\boldsymbol{P}_{t}^{\text{tea}}$. Overall, $\boldsymbol{P}_{v}^{\text{stu}}$ is continuously optimized by:
@@
[…]
@@
Datasets. Following the paradigm of benchmark settings in mainstream prompt tuning models (Zhou et al. 2022b; Khattak et al. 2023b; Li et al. 2024a), AugPT apply 11 recognition-related datasets with diverse domain distributions of data for evaluation on downstream tasks, containing ImageNet (Deng et al. 2009), Caltech101 (Fei-Fei, Fergus, and Perona 2004), OxfordPets (Parkhi et al. 2012), StanfordCars (Krause et al. 2013), Flowers102 (Nilsback and Zisserman 2008), Food101 (Bossard, Guillaumin, and Van Gool 2014), FGVCAircraft (Maji et al. 2013), SUN397 (Xiao et al. 2010), DTD (Cimpoi et al. 2014), EuroSAT (Helber et al. 2019) and UCF101 (Soomro 2012).
Baselines. We select 7 representative data-driven prompt tuning models as baselines, including foundational CoOp (Zhou et al. 2022b), PromptSRC (Khattak et al. 2023b), KAPT (Kan et al. 2023) that links structured knowledge bases, and typical LLM-guided methods: generating descriptive texts (HPT (Wang et al. 2024)), attribute augmentation (ArGue (Tian et al. 2024)), and fine-tuning optimization (LLaMP (Zheng et al. 2024)). Beyond these, we make a detailed comparison with PromptKD (Li et al. 2024a) (current SOTA), the distillation-based backbone model of AugPT, to quantify the improvement across various tasks. We compare more baselines in Supplementary Material B.1. Additionally, to further evaluate the adaptability of AugPT, in Supplementary Material B.5, we introduce UPL (Zang et al. 2022) to prove that AugPT is a plug-and-play approach for distillation-based prompt tuning frameworks.
[…]
@@
Considering the additional cost of collecting unlabeled images in realistic scenes, we introduce a more challenging Few-shot Base-to-New task for distillation-based prompt tuning models (PromptKD & AugPT) to evaluate under data scarcity conditions. Instead of using all unlabeled images, this task only samples few-shot unlabeled images per class to measure base-class fine-tuning and new-class generalization. As shown in Tab. 2, AugPT still outperforms PromptKD in HM performance across both 8-shot and 16-shot scenarios. Notably, the enhancements are more significant than full-dataset setting, particularly on datasets with extremely limited images (e.g., EuroSAT with $<$200 samples). These results suggest that the strategy of AugPT that exploits internal image features significantly enhances the model’s ability to fit target distributions in data-scarce scenarios, highlighting its practical value.
@@
To further validate the effectiveness of the augmentation strategy (ASA) proposed in AugPT, we replace ASA with other mainstream self-supervised augmentation methods for comparison. For fairness, the number of augmented views $N$ and all other sub-modules in AugPT maintain unchanged. As in Tab. 5, a key conclusion is that, compared to RandAugment (Cubuk et al. 2020) with fixed amplitude $A$, AugPT with dynamic amplitude sampled from $A^{(s)}\sim\text{U}(0,A_{\text{max}})$ performs more stable improvements, suggesting better generalization across diverse data distributions. Meanwhile, AugPT outperforms the common Random Resized Crop (Mao et al. 2022) in prompt tuning data processing and achieves performance better than FixMatch (Sohn et al. 2020).
@@
Figure 3: Relative to the backbone model, the HM performance improvement under different (a) augmented image number $N$, (b) step size $S$ and (c) filtering gate operator.

### v2

@@ Introduction
Bridging images and text, Vision-Language Models (VLMs, represented by CLIP (radford2021clip)) pre-trained on hyper-scale image-text pairs have demonstrated impressive cross-modal alignment and fusion capabilities (li2023blip; alayrac2022flamingo). To efficiently adapt pretrained VLMs for classification-related downstream tasks, CLIP-based prompt tuning is proposed as a parameter-efficient fine-tuning (PEFT) approach (xing2024peftsurvey). Freezing all parameters in CLIP encoders, prompt tuning introduces a set of lightweight learnable prompt vectors that replace hard prompts in raw inputs (zhou2022coop; zhou2022cocoop). As self-adaptive queries, prompt vectors guide the pretrained CLIP toward the target distribution through efficient fine-tuning.
@@
Given the rareness of annotated image-text pairs in target downstream tasks, extant prompt-tuning strategies (khattak2023maple; yao2023kgcoop; khattak2023promptsrc) are generally carried out under few-shot settings. Therefore, intuitively, introducing more relevant data to the tuning process can enhance recognition performance on target categories (base classes). Empirical evidence from numerous studies (zhou2022coop; khattak2023maple; khattak2023promptsrc) supports this conclusion. As illustrated in Fig. 1(a), with the increase of provided image-text pairs (shots), consistent improvement on fine-tuning performance can be observed on multiple backbone models. This indicates that the prompt tuning process is highly sensitive to the scale of data.
@@
Figure 1: (a) Impact of the amount of data (shots) per class on performance in prompt tuning models (khattak2023maple; yao2023kgcoop; khattak2023promptsrc); (b) Visualization of bad cases in existing self-supervised augmentation approaches. In comparison, AugPT obtains quality improvement as in (c) by filtering noisy samples, surpassing the current state-of-the-art PromptKD (li2024promptkd) as backbone model in (d) base-to-new tasks in both full and few-shot scenarios over 11 datasets.
@@
Motivated by this characteristic, numerous approaches explore optimizing prompt tuning at the data level. Current studies primarily focus on introducing external knowledge resources to boost base-class performance and generalization on out-of-distribution categories (new classes), containing pre-built hard prompt templates (yao2023kgcoop; yu2023taskres) or structured knowledge bases (kan2023kapt; zhang2024cocole), and knowledge generated by Large Language Models (LLMs) (wang2024hpt; tian2024argue). Despite the effectiveness, these strategies require additional data collection and organization costs. In addition, reliance on task-specific external knowledge also limits the generality and transferability for fine-tuning the models. Latest research (li2024promptkd; wu2024caspl; li2025dpc) further proposes methods for data amplification utilizing internal knowledge of the model. Typically guided by knowledge distillation, these approaches first fine-tune a prompt tuning backbone under few-shot settings (khattak2023promptsrc) as teacher model, then guide it to execute knowledge completion for available unlabeled images for further fine-tuning a student model. Although distillation-based method achieve state-of-the-art (SOTA) in both base-class accuracy and new-class generalization, there are still notable limitations: (1) Performance gains still depend on the expansion of unlabeled image library, which also requires extra data collection costs; (2) Extracted image features are only applied for aligning teacher and student, signifying that raw images are not adequately exploited.
@@
It is worth noting that traditional self-supervised augmentation strategies (cubuk2020randaugment; devries2017cutout; mao2022randomresizedcrop) face a key limitation: they lack an error correction mechanism, which may introduce noisy samples that discard key visual features, as illustrated in Fig. 1(b). Although this limitation can be addressed with additional supervision (sohn2020fixmatch), to maintain parameter efficiency of prompt tuning, we consider strategies that do not require external supervision. As an effective solution, AugPT introduces a novel Consensus Test filtering gate, leveraging a previously overlooked property of pre-tuned prompt tuning models: logit-level consistency reflects semantic similarity across images (details are in Method section). Concretely, we reuse the frozen teacher model (li2024promptkd) to perform online inference over both raw and augmented images, and adopt Top-1 frequency consistency as consensus to build the view filtering strategy, ensuring that core semantic features of raw image are preserved in augmented views (Fig. 1(c)). Without the need of external modules, this design injects a self-corrective mechanism into augmentation, thereby enhancing cross-modal alignment.
Experiments across 11 datasets reveal that AugPT exceeds PromptKD (li2024promptkd) (current SOTA) as backbone model on base-class performance, new-class generalization, and cross-dataset transfer, without using additional knowledge or external data. Our main contributions are:
@@
Prompt Tuning in VLMs. Although CLIP (radford2021clip)-based VLMs has remarkable performance on cross-modal tasks, the Transformer-based encoders (dosovitskiy2020vit) with large parameters make it challenging on full fine-tuning for adapting downstream tasks. Therefore, prompt tuning is proposed as a parameter-efficient strategy (xing2024peftsurvey; han2024peftsurveyindpc). It does not rely on hard prompts (e.g., “A photo of a [CLASS]”), but introduces a set of learnable vectors as substitutes. During fine-tuning, all parameters of encoders are frozen, and the prompt vectors used as queries are continuously optimized to fit the distribution of target tasks. Existing research explores various formats of prompts, containing text re-encoding (zhou2022coop; zhu2023prograd; tian2024argue), concatenation with image features (jia2022vpt), and joint visual-text encoding (khattak2023maple; khattak2023promptsrc; li2025mao; guo2025mmrl). Since CLIP-based prompt tuning is typically conducted under few-shot scenarios, applying data amplification strategies like AugPT to enhance performance is intuitive, which is supported by the results in Fig. 1(a).
Figure 2: Framework of AugPT. In (a) fine-tuning stage, AugPT applies Adaptive Augmentation to raw unlabeled images $I^{\prime}$, then reuses the teacher model from distillation-based backbone to discard noisy samples in the augmented image set $\mathcal{D}(I^{\prime})$ via Top-1 Consensus-based Filtering Gate, followed by fine-tuning through Optimized Prompt Distillation by fitting the logits of augmented data between teacher and student. Tuned student image prompts $\boldsymbol{P}_{v}^{\text{stu}}$ and projection layer $\text{Proj}(\cdot)$ are applied in (b) inference stage to align student visual features with teacher text embeddings $g^{\text{tea}}({\boldsymbol{P}_{t}^{\text{tea}}})$.
Data-Driven Optimization for Prompt Tuning. Given the limitation of image-text pairs for fine-tuning in CLIP-based prompt tuning, a variety of studies discover data-level optimization strategies. Mainstream approaches consider introducing external knowledge to expand data or constrain the optimizing flow, including elaborately designed hard prompt templates as supervision (yao2023kgcoop; khattak2023promptsrc), or structured databases for appending additional knowledge (kan2023kapt; zhang2024cocole). More recently, LLMs are employed as external sources to generate descriptions (pratt2023cupl; wang2024hpt; khattak2025protext), auxiliary attributes (tian2024argue; li2024atprompt; zhu2024awt), or guide fine-tuning (zheng2024llamp). Since the above methods demand extra constructing cost for external knowledge bases, the latest works (mistretta2024kdpl; li2024promptkd) propose distillation-based paradigms using internal knowledge + external data. Specifically, a teacher model is first fine-tuned on few-shot image-text pairs and then guides a student model by knowledge completion to external unlabeled images. However, further optimizing such models still requires scaling up external data pool, signifying additional collection overhead. In contrast, AugPT proposes an optimized distillation-based strategy that relies only on internal knowledge + internal augmentation based on existing raw data.
@@
The framework of AugPT is illustrated in Fig. 2. AugPT adopts the infrastructure of an up-to-date distillation-based prompt tuning model (taking SOTA PromptKD (li2024promptkd) as a paradigm) as its backbone. Existing unlabeled images from the backbone are sequentially processed through an Adaptive Self-supervised Augmentation module and a Consensus-based Filtering Gate. The pipeline is then optimized for further exploiting latent features within the raw data during fine-tuning.
@@
Herein, $\text{Proj}(\cdot)$ is a learnable MLP-based projection layer, aligning the shape of visual embeddings of student with the text features of teacher model for modal interaction.
@@
To avoid introducing additional learnable layers or external knowledge, thereby minimizing the cost of data collection and computational overhead, AugPT adopts parameter-free image augmentation based on self-supervised strategy. Replacing the naive Random Resized Crop (mao2022randomresizedcrop) applied in standard prompt tuning, based on the RandAugment (cubuk2020randaugment) methodology, AugPT performs $N$ augmentations on each raw unlabeled image $I^{\prime}$. For each augmentation, the model samples an operation from a pre-designed set with $Q$ candidates of mainstream transformation policies $\mathcal{Y}=\{\gamma_{1},\dots,\gamma_{Q}\}$ (details are in Supplementary Material A.3). Guided by a searching space defined by step size $S$ and amplitude $A$, augmentation operations $\gamma^{(s)}\sim\text{Uniform}(\mathcal{Y})$ are independently and identically sampled at each step $s\in[S]$ to form a policy group $\mathcal{G}$:
@@
Based on the filtered image set $\mathcal{I}_{\text{acc}}$ and the consensus category $\theta^{\star}$, AugPT executes further optimization following the fine-tuning workflow of distillation-based prompt tuning backbone models (li2024promptkd). Specifically, under the constraint of $\theta^{\star}$, AugPT samples the teacher logits $\hat{\mathcal{\ell}}^{\text{tea}}_{p}$ corresponding to each accepted image $\hat{I}_{p}\in\mathcal{I}_{\text{acc}}$ to organize a set $\mathcal{T}^{\text{tea}}_{\text{acc}}$ in a one-to-one correspondence manner:
@@
Next, AugPT constructs a student model identical to the backbone (which a frozen CLIP-ViT-B/16 image encoder $f^{\text{stu}}(\cdot)$ and learnable student image prompts $\boldsymbol{P}_{v}^{\text{stu}}$). Similarly, KL divergence is utilized to align the distributions between teacher logits $\hat{\mathcal{\ell}}^{\text{tea}}_{p}$ from the filtered set and student logits $)]\hat{\mathcal{\ell}}^{\text{stu}}_{p}=sim[g^{\text{tea}}({\boldsymbol{P}_{t}^{\text{tea}}}),\text{Proj}(f^{\text{stu}}({\boldsymbol{P}_{v}^{\text{stu}}},\hat{I_{p}}$))], which are obtained by interacting the learnable student visual inputs $(\boldsymbol{P}_{v}^{\text{stu}},\hat{I}_{p})$ with the frozen teacher text prompts $\boldsymbol{P}_{t}^{\text{tea}}$. Overall, $\boldsymbol{P}_{v}^{\text{stu}}$ is continuously optimized by:
@@
Herein, $\text{Proj}(\cdot)$ refers to an MLP-based learnable projection layer that is identical to the backbone model (li2024promptkd), which aligns the student’s ViT-B/16-based visual features ($dim=512$) with the teacher’s ViT-L/14-based text features ($dim=768$). Aforementioned fine-tuning process promotes AugPT to optimize cross-modal alignment across all filtered augmented views, more effectively exploiting the latent features in the raw data.
@@
Datasets. Following the paradigm of benchmark settings in mainstream prompt tuning models (zhou2022coop; khattak2023promptsrc; li2024promptkd), AugPT apply 11 recognition-related datasets with diverse domain distributions of data for evaluation on downstream tasks, containing ImageNet (deng2009imagenet), Caltech101 (fei2004caltech), OxfordPets (parkhi2012pets), StanfordCars (krause2013cars), Flowers102 (nilsback2008flowers), Food101 (bossard2014food), FGVCAircraft (maji2013aircraft), SUN397 (xiao2010sun), DTD (cimpoi2014dtd), EuroSAT (helber2019eurosat) and UCF101 (soomro2012ucf101).
Baselines. We select 7 representative data-driven prompt tuning models as baselines, including foundational CoOp (zhou2022coop), PromptSRC (khattak2023promptsrc), KAPT (kan2023kapt) that links structured knowledge bases, and typical LLM-guided methods: generating descriptive texts (HPT (wang2024hpt)), attribute augmentation (ArGue (tian2024argue)), and fine-tuning optimization (LLaMP (zheng2024llamp)). Beyond these, we make a detailed comparison with PromptKD (li2024promptkd) (current SOTA), the distillation-based backbone model of AugPT, to quantify the improvement across various tasks. We compare more baselines in Supplementary Material B.1. Additionally, to further evaluate the adaptability of AugPT, in Supplementary Material B.5, we introduce UPL (zang2022upt) to prove that AugPT is a plug-and-play approach for distillation-based prompt tuning frameworks.
Implementation Details. We initialize backbone model of AugPT using the optimal parameters obtained through prior exploration by PromptKD, including the pre-tuned teacher model with ViT-L/14 encoders, as well as a student model with a ViT-B/16 encoder $f^{\text{stu}}(\cdot)$ and learnable image prompts $\boldsymbol{P}_{v}^{\text{stu}}$. Identical to PromptKD, $\text{Proj}()$ layer consists of 2 layers. For internal data augmentation, based on ablations, we set the number of augmented images to $N=5$ and the augmenting step to $S=2$. Further details and hyperparameters are enumerated in Supplementary Material A.
@@
Considering the additional cost of collecting unlabeled images in realistic scenes, we introduce a more challenging Few-shot Base-to-New task for distillation-based prompt tuning models (PromptKD & AugPT) to evaluate under data scarcity conditions. Instead of using all unlabeled images, this task only samples few-shot unlabeled images per class to measure base-class fine-tuning and new-class generalization. As shown in Tab. 2, AugPT still outperforms PromptKD in HM performance across both 8-shot and 16-shot scenarios. Notably, the enhancements are more significant than full-dataset setting, particularly on datasets with extremely limited images (e.g., EuroSAT with <<200 samples). These results suggest that the strategy of AugPT that exploits internal image features significantly enhances the model’s ability to fit target distributions in data-scarce scenarios, highlighting its practical value.
@@
To further validate the effectiveness of the augmentation strategy (ASA) proposed in AugPT, we replace ASA with other mainstream self-supervised augmentation methods for comparison. For fairness, the number of augmented views $N$ and all other sub-modules in AugPT maintain unchanged. As in Tab. 5, a key conclusion is that, compared to RandAugment (cubuk2020randaugment) with fixed amplitude $A$, AugPT with dynamic amplitude sampled from $A^{(s)}\sim\text{U}(0,A_{\text{max}})$ performs more stable improvements, suggesting better generalization across diverse data distributions. Meanwhile, AugPT outperforms the common Random Resized Crop (mao2022randomresizedcrop) in prompt tuning data processing and achieves performance better than FixMatch (sohn2020fixmatch).
@@
Figure 3: Relative to the backbone model, the HM performance improvement under different (a) augmented image number NN, (b) step size $S$ and (c) filtering gate operator.